# The Nonlinear Filter Model of Stream Cipher Redivivus

### v6

@@ 4.1 Effect of Shifts on Tap Positions
Suppose $\kappa$, $L\geq 2\kappa$, and $m$ are given. We need to determine $i_{1},\ldots,i_{m}$ and $j_{1},\ldots,j_{m}$, or equivalently, to determine ${\sf pos}$. (Recall that the tap position for $W$ is fixed to be $L-2\kappa$.) The leftmost $\kappa$ bits of ${\sf pos}$ is the string ${\sf posX}$. The value of $\nu$ is determined by ${\sf posX}$, while the value of $\delta$ is determined by ${\sf pos}$. We would like to minimise $\nu$ as well as $\delta$. It is difficult to simultaneously minimise both. From (20) (provided later), the value of $\delta$ is one of the factors relevant to protection against certain kinds of state guessing attacks. As explained in Remark 2 (also provided later), it is not essential to minimise $\delta$. The parameter $\nu$, on the other hand, determines the size of the maximum overlap between two inputs to the majority function corresponding to two different keystream bits. Minimising $\nu$ improves resistance to possible differential attacks (see Section 4.4). Since it is not essential to minimise $\delta$, we chose to focus on minimising $\nu$. The task of choosing ${\sf posX}$ such that $\nu$ is minimised is a combinatorial optimisation problem. There does not seem to be any good way to obtain ${\sf posX}$ such that the corresponding $\nu$ is guaranteed to be the minimum possible value. Instead, we adopted the following procedure. Given $\kappa$, $L$ and $m$, make 10000 choices of ${\sf pos}$ satisfying the constraints on $i_{1},\ldots,i_{m}$ and $j_{1},\ldots,j_{m}$, and for each choice extract ${\sf posX}$ and then compute $\nu$ from ${\sf posX}$. Finally return the ${\sf pos}$ for which the least value of $\nu$ is obtained. Then, for this ${\sf pos}$ compute the value of $\delta$. Later we provide the values of ${\sf pos}$, $\nu$ and $\delta$ for the concrete proposals that we make.
@@ 4.2 Security Analysis
We consider two basic attack parameters, namely the number $N$ of keystream bits that is required for a successful attack, and the time complexity $T$ of the attack. To ensure security at level $\kappa$ any attack should require $T>2^{\kappa}$. The condition $L\geq 2\kappa$ ensures that the size of the state is at least twice the size of the key. This prevents certain (theoretical) time/memory trade-off attacks [34]. We assume that at most $2^{B}$ keystream bits are generated from a single key and IV pair. So $N$ is at most $2^{B}$, as otherwise the attack cannot be mounted. Further, a basic condition is that $N\leq T$, i.e. some operation is performed on each keystream bit that is required. So for $\kappa$-bit security using $2^{B}$ keystream bits, we have $B\leq\kappa$.
@@
The degree of $({\sf Maj},{\sf rev})\mbox{-}{\sf MM}_{2m+1}$ is $d=2^{\lfloor\log_{2}m\rfloor}$. If $L$ is a prime, then the linear complexity of the generated keystream sequence is (${L\choose d}$ (see [49]) and so at least these many keystream bits are required to determine the linear complexity. So if
@@
Suppose at some point of keystream generation, the state is $\mathbf{s}^{(0)}=(s_{L-1},\ldots,s_{0})$ and the keystream bit generated from $\mathbf{s}^{(0)}$ is $w_{0}$. From the definition of the filtering function, the bit $w_{0}$ is generated from the bits $(s_{L-1},\ldots,s_{L-2\kappa})$, and we can write $w_{0}=s_{L-2\kappa}\oplus g(s_{L-1},\ldots,s_{L-2\kappa+1})$, where $g$ is some function (defined from ${\sf MM}_{2m}$ and is non-degenerate on $2m$ variables). Recall that the tap position for $X_{1}$ is $L-1$ and the tap position for $Y_{m}$ is $L-2\kappa+1$. The gap between the tap position of $X_{1}$ and the tap position of $Y_{m}$ (including both of these tap positions) is $\chi=2\kappa-1$. Note that $s_{L-2\kappa}=w_{0}\oplus g(s_{L-1},\ldots,s_{L-2\kappa+1})$. Extending this equation in the backward direction, we obtain the following equations.
@@
The sequence $w_{0},w_{-1},w_{-2},\ldots$ is known. If we guess the values of $s_{L-1},\ldots,s_{L-2\kappa+1}$ (a total of $\chi$ bits), then we obtain $s_{L-2\kappa}$ from the first equation, using the obtained value of $s_{L-2\kappa}$ in the second equation, we obtain the value of $s_{L-2\kappa-1}$, using the obtained value of $s_{L-2\kappa-1}$ in the third equation, we obtain the value of $s_{L-2\kappa-2}$, and so on. In other words, by guessing the values of $\chi$ of the state bits, we can obtain values of $L-\chi$ previous state bits, giving us the values of all $L$ bits of the state at a point $L-\chi$ steps in the backward direction. Once the complete state is known, the original guess can be verified by generating the keystream in the forward direction and matching with the obtained keystream. The method requires $2^{\chi}$ guesses, and for each guess about $Lm$ operations are required. By our choice of the tap positions for $\mathbf{X}$ and $\mathbf{Y}$, the value of $\chi$ is $12\kappa-$ 1. So the number of guesses required to successfully mount the attack is far greater than $2^{\kappa}$. This proves the resistance of the stream cipher to the inversion attack. Note that the inversion attack is prevented due to $\chi$ being sufficiently large. This was one of the countermeasures to the inversion attack that was already proposed in [27].
@@
Let as above the state of the LFSR at some point be $\mathbf{s}^{(0)}$ and the subsequent states of the LFSR be denoted as $\mathbf{s}^{(1)},\mathbf{s}^{(2)},\ldots,\mathbf{s}^{(c-1)}$. For $i=0,\ldots,c-1$, let $\mathbf{t}^{(i)}={\sf proj}(\mathbf{s}^{(i)})$ and $w_{i}=f_{2m+1}(\mathbf{t}^{(i)})$. Recall from the definition of $\delta$ that for $10\leq i<j\leq c-$ 1, $\mathbf{t}^{(i)}$ and $\mathbf{t}^{(j)}$ have an overlap of at most $\delta$ bits. So in the $c$ $n$-bit strings $\mathbf{t}^{(0)},\ldots,\mathbf{t}^{(c-1)}$ there are at least $nc-\delta c(c-1)/2$ unknown bits and hence there are at least $2^{nc-\delta c(c-1)/2}$ possibilities for these $c$ strings. Since $f_{2m+1}$ is balanced, knowledge of $w_{0}$ reduces the number of possibilities for $\mathbf{t}^{(0)}$ by a factor of half, and more generally, the knowledge of $w_{0},\ldots,w_{c-1}$ reduces the number of possibilities for $\mathbf{t}^{(0)},\ldots,\mathbf{t}^{(c-1)}$ by a factor of $2^{c}$. So at least $2^{(n-1)c-\delta c(c-1)/2}$ possibilities remain for the $c$ strings $\mathbf{t}^{(0)},\ldots,\mathbf{t}^{(c-1)}$. For each of these possibilities, we obtain a system of $L$ equations in $L$ variables, the solution of which provides the value of initial state of the LFSR. The correctness of the obtained value of the state can be determined by generating the keystream from the obtained value and matching with the actual keystream. Solving the system of $L$ linear equations requires $L^{3}$ arithmetic operations, and generating the keystream from the obtained state require about $Ln$ operations. So the complexity of the attack is at least $2^{(n-1)c-\delta c(c-1)/2}(L^{3}+Ln)$ operations. Suppose the following condition holds.
$2^{(n-1)c-\delta c(c-1)/2}(L^{3}+Ln)$ $>$ $2^{\kappa}\mbox{ where $c$ is the least positive integer such that $nc\geq L$.}$ (20)
@@
Applying the attack requires knowing the pre-image sets of 00 and $1$ of the function $f_{2m+1}$. These need to be stored separately, and depending upon the value of $w_{i}$, the appropriate pre-image set is to be used. So the storage required is $2^{2m+1}$. For the concrete proposals that we put forward later, the value of $n=2m+1$ is quite close to that of the security level $\kappa$ (see Table 1). As a result, the memory requirement will be prohitively high to apply the attack.
The complexity of the above attack depends on the value of $\delta$, which is the maximum overlap of tap positions between shifts of the state vector. The recommendations for tap positions in the literature [27, 28, 33, 47, 58] aim to reduce this overlap. A commonly used recommendation is full positive difference set, i.e. the absolute values of the differences between the tap positions should be distinct. For a filtering function with $n$ variables, this recommendation results in the gap between the first and the last tap positions to be more than $21+2+\cdots+(n-1)=n(n-1)/$ 2 (since the successive differences must be distinct). So the size of the LFSR is at least quadratic in $n$. From efficiency considerations, this forces the value of $n$ to be small. By not following this recommendation, we have done away with the condition that $L$ must be at least quadratic in $n$. Instead, by considering the fundamental requirement behind state guessing attacks, we identified (20) as the condition to resist such attacks. This allows us to choose $n$ to be quite large and close to $\kappa$. Of course, this is possible due to the fact that we have an efficient method for implementing the filtering functions on such large values of $n$.
@@
Type of attack. For attacks based on low-weight parity-check equations [45, 13], the number of keystream bits required is about $N=(2\varepsilon)^{-2(d-2)/(d-1)}\cdot 2^{L/(d-1)}$, the pre-computation step requires about $N^{d-2}/(d-2)!$ operations, and the (online) time for decoding is about $(2\varepsilon)^{-2d(d-2)/(d-1)}\cdot 2^{L/(d-1)}$, where $d\geq 3$ is the number of non-zero terms in (some multiple of the) LFSR connection polynomial, and $\varepsilon$ is the linear bias.
Evaluation against $\mathcal{S}(L,m)$. We have $\varepsilon=2^{-m-1}$ and so $N=2^{(2(m+2)(d-2)+L)/(d-1)}$. We set $N=2^{B}$ and so $(d-1)(B-2(m+2))=L-2(m+2)$. If $B=2(m+2)$, then $L=2(m+2)$. We choose $L$ and $m$ to ensure that $L>2(m+2)$, and so $B\neq 2(m+2)$. In this case, we solve for $d$ to obtain $d=1+(L-2(m+2))/(B-2(m+2))$. If $B<2(m+2)$, then $d<1$ which violates the condition $d\geq 3$ and the attack does not work. So let us consider $B>2(m+2)$. In this case, substituting the obtained expression for $d$ in the expression for the decoding time, we find the decoding time to be $2^{(2(m+2)L+B^{2})/(B-2(m+2))}$. So the following condition ensures $\kappa$-bit security.
$L>(2m+2)\mbox{ and either }B<2(m+2)\mbox{ or }\kappa<\frac{2(m+2)L+B^{2}}{B-2(m+2)}.$ (21)
@@
Evaluation against $\mathcal{S}(L,m)$. Again we have $\varepsilon=2^{-m-1}$. Setting $N=2^{B}$, we obtain $k=L+2w(m+1)-wB$. Setting $T$ to be the time complexity of the decoding step, we obtain $\log_{2}T=L+2w(2m+3)-wB$. Since $L\geq 2\kappa$, ensuring $2(2m+3)\geq B$ is (more than) sufficient to ensure $\kappa$-bit security, irrespective of the values of $w$ and $k$. We record this condition as follows.
@@
Evaluation against $\mathcal{S}(L,m)$. The size of the support of the Walsh transform of ${\sf MM}_{2m+1}$ is equal to the size of the support of the Walsh transform of ${\sf MM}_{2m}$. Since ${\sf MM}_{2m}$ is bent, it follows that the size of the support of ${\sf MM}_{2m}$ is $2^{2m}$. So $F=2^{2m}$. Setting $N=2^{B}$, we obtain $k=L-wB$. Substituting $k$ in the expression for $T$, we obtain $\log_{2}T=L+w(2m-B)$. Since $L\geq 2\kappa$ and $w\geq 2$, we have $\log_{2}T=L+w(2m-B)\geq 2\kappa+2(2m-B)$. So if $2\kappa+2(2m-B)>\kappa$, or equivalently, $\kappa+4m>2B$, then $\kappa$-bit security is achieved against the attack. We record this condition as follows.
@@
For a filtering function $f$ having ${\sf AI}(f)=a$, an algebraic attack requires about $N=\sum_{i=0}^{a}{L\choose i}$ keystream bits and has time complexity $T$ to be about $N^{\omega}$, where $\omega$ is the exponent of matrix multiplication (see Section 3.1.5 of [14]). We assume $\omega=2.8$. For our proposal, the filtering function is $f_{2m+1}$ whose algebraic resistance is the same as that of $({\sf Maj},{\sf rev})\mbox{-}{\sf MM}_{2m+1})$. Let $a={\sf(}f_{2m+1})={\sf AI}(({\sf Maj},{\sf rev})\mbox{-}{\sf MM}_{2m+1})$. From Proposition 4, we assume that $a=\lceil m/2\rceil$, i.e. the actual algebraic immunity of $f_{2m+1}$ is equal to the lower bound on the algebraic immunity. Let
@@
For any $(e,d)$ for which there are functions $g$ and $h$ of degrees $e$ and $d$ respectively such that $gf_{2m+1}=h$, the Berlekamp-Massey step in a fast algebraic attack on the filter generator model has time complexity $\mathcal{O}(ED\log D)$, where $E=\sum_{i=0}^{e}{L\choose i}$ and $D=\sum_{i=0}^{d}{L\choose i}$ (see [31] and Section 3.1.5 of [14]). This complexity dominates the overall time complexity of a fast algebraic attack. Ignoring the logarithm term, we take $T$ to be equal to $ED$. The number $N$ of keystream bits required is about $2E$. The maximum value of $e$ is one less than the algebraic immunity $a$. From Proposition 4, we again assume as above that $a=\lceil m/2\rceil$. Recall that for any Boolean function, its fast algebraic immunity is at least one more than its algebraic immunity, i.e. ${\sf FAI}(f_{2m+1})\geq a+1$. For any functions $g$ and $h$ of degrees $e$ and $d$ respectively such that $gf_{2m+1}=h$, we have $e+d\geq{\sf FAI}(f_{2m+1})\geq a+1$, and we assume that the lower bound is the actual value of $e+d$, and so $d=a+1-e$. Define

### v7

@@ 4.1 Effect of Shifts on Tap Positions
Suppose $\kappa$, $L\geq 2\kappa$, and $m$ are given. We need to determine $i_{1},\ldots,i_{m}$ and $j_{1},\ldots,j_{m}$, or equivalently, to determine $\sf pos$}. (Recall that the tap position for $W$ is fixed to be $L-2\kappa$.) The leftmost $\kappa$ bits of $\sf pos$} is the string $\sf posX$}. The value of $\nu$ is determined by $\sf posX$}, while the value of $\delta$ is determined by $\sf pos$}. We would like to minimise $\nu$ as well as $\delta$. It is difficult to simultaneously minimise both. From (27) (provided later), the value of $\delta$ is one of the factors relevant to protection against certain kinds of state guessing attacks. As explained in Remark 2 (also provided later), it is not essential to minimise $\delta$. The parameter $\nu$, on the other hand, determines the size of the maximum overlap between two inputs to the majority function corresponding to two different keystream bits. Minimising $\nu$ improves resistance to possible differential attacks (see Section 4.4). Since it is not essential to minimise $\delta$, we chose to focus on minimising $\nu$. The task of choosing $\sf posX$} such that $\nu$ is minimised is a combinatorial optimisation problem. There does not seem to be any good way to obtain $\sf posX$} such that the corresponding $\nu$ is guaranteed to be the minimum possible value. Instead, we adopted the following procedure. Given $\kappa$, $L$ and $m$, we randomly generated the array $\sf pos$} in the following manner. Set ${\sf pos}[L-2\kappa]=1$ (the tap position for the variable $W$); set $i_{1}=0$, randomly select a subset $\{i_{2},\ldots,i_{m}\}$ of size m1m-1 from $\{1,\ldots,\kappa-1\}$ and set ${\sf pos}[L-1-i_{p}]=1$, for $p=1,\ldots,m$ (these provide the tap positions for the variables $X_{1},\ldots,X_{m}$); set $j_{m}=2\kappa-2$, randomly select a subset $\{j_{1},\ldots,j_{m-1}\}$ from $\{\kappa,\ldots,2\kappa-3\}$ and set ${\sf pos}[L-1-j_{p^{\prime}}]=1$ (these provide the tap positions for the variables $Y_{1},\ldots,Y_{m}$). From the obtained $\sf pos$}, we extract $\sf posX$} and then we compute $\nu$ from $\sf posX$}. So for each random choice of $\sf pos$}, we obtain a value of $\nu$. The random procedure is repeated 10000 times, providing 10000 values of $\nu$. Finally we return the $\sf pos$} for which the least value of $\nu$ is obtained. Then, for this $\sf pos$} compute the value of $\delta$. Through experimentation we found that repeating the random procedure more than 10000 times does not result in lower values of $\nu$. Later we provide the values of $\sf pos$}, $\nu$ and $\delta$ for the concrete proposals that we make.
@@ 4.2 Security Analysis
We consider the security of the stream cipher. For the analysis of security, we consider key recovery attacks, i.e. the goal of the adversary is to recover the secret key. Since both the next state map $\sf NS$} (see (8)) of the LFSR and the initialisation round function $\sf IR$} (see (18)) are invertible, it follows that recovering the state of the stream cipher at any point of time essentially amounts to recovering the secret key, i.e. state recovery implies efficient key recovery.
@@
We consider two basic attack parameters, namely the number $N$ of keystream bits that is required for a successful attack, and the time complexity $T$ of the attack. By the time complexity we mean the number of bit operations required to carry out the attack. To ensure security at level $\kappa$ any attack should require $T>2^{\kappa}$. The condition $L\geq 2\kappa$ ensures that the size of the state is at least twice the size of the key. This prevents certain time/memory trade-off attacks.
Suppose $B$ is a parameter such that at most $2^{B}$ keystream bits are generated from a single key and IV pair. So $N$ is at most $2^{B}$, as otherwise the attack cannot be mounted. Further, a basic condition is that $N\leq T$, i.e. some operation is performed on each keystream bit that is required. So for $\kappa$-bit security using $2^{B}$ keystream bits, we have $B\leq\kappa$. Recall that for our stream cipher proposal we have fixed $B=64$. For the description of the attacks we proceed without this restriction.
@@
The degree of $({\sf Maj},{\sf rev})\mbox{-}{\sf MM}_{2m+1}$ is $d=2^{\lfloor\log_{2}m\rfloor}$. If $L$ is a prime, then the linear complexity of the generated keystream sequence is $L\choose d$} (see [49]) and so at least this many keystream bits are required to determine the linear complexity. So if
@@
Suppose at some point of keystream generation, the state is $\mathbf{s}^{(0)}=(s_{L-1},\ldots,s_{0})$ and the keystream bit generated from $\mathbf{s}^{(0)}$ is $w_{0}$. From the definition of the filtering function, the bit $w_{0}$ is generated from the bits $(s_{L-1},\ldots,s_{L-2\kappa})$, and we can write $w_{0}=s_{L-2\kappa}\oplus g(s_{L-1},\ldots,s_{L-2\kappa+1})$, where $g$ is some function (defined from ${\sf MM}_{2m}$ and is non-degenerate on $2m$ variables). Recall that the tap position for $X_{1}$ is L1L-1 and the tap position for $Y_{m}$ is $L-2\kappa+1$. The gap between the tap position of $X_{1}$ and the tap position of $Y_{m}$ (including both of these tap positions) is $\chi=2\kappa-1$. Note that $s_{L-2\kappa}=w_{0}\oplus g(s_{L-1},\ldots,s_{L-2\kappa+1})$. Extending this equation in the backward direction, we obtain the following equations.
@@
The sequence $w_{0},w_{-1},w_{-2},\ldots$ is known. If we guess the values of $s_{L-1},\ldots,s_{L-2\kappa+1}$ (a total of $\chi$ bits), then we obtain $s_{L-2\kappa}$ from the first equation, using the obtained value of $s_{L-2\kappa}$ in the second equation, we obtain the value of $s_{L-2\kappa-1}$, using the obtained value of $s_{L-2\kappa-1}$ in the third equation, we obtain the value of $s_{L-2\kappa-2}$, and so on. In other words, by guessing the values of $\chi$ of the state bits, we can obtain values of $L-\chi$ previous state bits, giving us the values of all $L$ bits of the state at a point $L-\chi$ steps in the backward direction. Once the complete state is known, the original guess can be verified by generating the keystream in the forward direction and matching with the obtained keystream. The method requires $2^{\chi}$ guesses, and for each guess about $Lm$ bit operations are required. By our choice of the tap positions for $\mathbf{X}$ and $\mathbf{Y}$, the value of $\chi$ is $2\kappa-1$. So the number of guesses required to successfully mount the attack is far greater than $2^{\kappa}$. This proves the resistance of the stream cipher to the inversion attack. Note that the inversion attack is prevented due to $\chi$ being sufficiently large. This was one of the countermeasures to the inversion attack that was already proposed in [27].
@@
Let as above the state of the LFSR at some point be $\mathbf{s}^{(0)}$ and the subsequent states of the LFSR be denoted as $\mathbf{s}^{(1)},\mathbf{s}^{(2)},\ldots,\mathbf{s}^{(c-1)}$. For $i=0,\ldots,c-1$, let $\mathbf{t}^{(i)}={\sf proj}(\mathbf{s}^{(i)})$ and $w_{i}=f_{2m+1}(\mathbf{t}^{(i)})$. Recall from the definition of $\delta$ that for $0\leq i<j\leq c-1$, $\mathbf{t}^{(i)}$ and $\mathbf{t}^{(j)}$ have an overlap of at most $\delta$ bits. So in the $c$ $n$-bit strings $\mathbf{t}^{(0)},\ldots,\mathbf{t}^{(c-1)}$ there are at least $nc-\delta c(c-1)/2$ unknown bits and hence there are at least $2^{nc-\delta c(c-1)/2}$ possibilities for these $c$ strings. Since $f_{2m+1}$ is balanced, knowledge of $w_{0}$ reduces the number of possibilities for $\mathbf{t}^{(0)}$ by a factor of half, and more generally, the knowledge of $w_{0},\ldots,w_{c-1}$ reduces the number of possibilities for $\mathbf{t}^{(0)},\ldots,\mathbf{t}^{(c-1)}$ by a factor of $2^{c}$. So at least $2^{(n-1)c-\delta c(c-1)/2}$ possibilities remain for the $c$ strings $\mathbf{t}^{(0)},\ldots,\mathbf{t}^{(c-1)}$. For each of these possibilities, we obtain a system of $L$ equations in $L$ variables, the solution of which provides the value of initial state of the LFSR. The correctness of the obtained value of the state can be determined by generating the keystream from the obtained value and matching with the actual keystream. Solving the system of $L$ linear equations requires $L^{3}$ bit operations, and generating the keystream from the obtained state require about $Ln$ bit operations. So the complexity of the attack is at least $2^{(n-1)c-\delta c(c-1)/2}(L^{3}+Ln)$ operations. Suppose the following condition holds.
[…]
@@
Applying the attack requires knowing the pre-image sets of 0 and $1$ of the function $f_{2m+1}$. These need to be stored separately, and depending upon the value of $w_{i}$, the appropriate pre-image set is to be used. So the storage required is $2^{2m+1}$. For the concrete proposals that we put forward later, the value of $n=2m+1$ is quite close to that of the security level $\kappa$ (see Table 1). As a result, the memory requirement will be prohitively high to apply the attack.
The complexity of the above attack depends on the value of $\delta$, which is the maximum overlap of tap positions between shifts of the state vector. The recommendations for tap positions in the literature [27, 28, 33, 47, 58] aim to reduce this overlap. A commonly used recommendation is full positive difference set, i.e. the absolute values of the differences between the tap positions should be distinct. For a filtering function with $n$ variables, this recommendation results in the gap between the first and the last tap positions to be more than $1+2+\cdots+(n-1)=n(n-1)/2$ (since the successive differences must be distinct). So the size of the LFSR is at least quadratic in $n$. From efficiency considerations, this forces the value of $n$ to be small. By not following this recommendation, we have done away with the condition that $L$ must be at least quadratic in $n$. Instead, by considering the fundamental requirement behind state guessing attacks, we identified (27) as the condition to resist such attacks. This allows us to choose $n$ to be quite large and close to $\kappa$. Of course, this is possible due to the fact that we have an efficient method for implementing the filtering functions on such large values of $n$.
@@
Type of attack. For attacks based on low-weight parity-check equations [45, 13], the number of keystream bits required is about $N=(2\varepsilon)^{-2(d-2)/(d-1)}\cdot 2^{L/(d-1)}$, the pre-computation step requires about $N^{d-2}/(d-2)!$ operations, and the (online) time for decoding is about $(2\varepsilon)^{-2d(d-2)/(d-1)}\cdot 2^{L/(d-1)}$, where $d\geq 3$ is the number of non-zero terms in (some multiple of the) LFSR connection polynomial, and $\varepsilon$ is the linear bias. Note that for this attack, the parameter $d$ is not the algebraic degree of the filtering function.
Evaluation against $\mathcal{S}(L,m)$. We have $\varepsilon=2^{-m-1}$ and so $N=2^{2m(d-2)+L)/(d-1)}$. We set $N=2^{B}$ and so (d1)(B2m)=L2m(d-1)(B-2m)=L-2m. If $B=2m$, then $L=2m$. We choose $L$ and $m$ to ensure that $L>2m$, and so $B\neq 2m$. In this case, we solve for $d$ to obtain d=1+(L2m)/(B2m)d=1+(L-2m)/(B-2m). If $B<2m$, then $d<1$ which violates the condition $d\geq 3$ and the attack does not work. So let us consider $B>2m$. In this case, substituting the obtained expression for $d$ in the expression for the decoding time, we find the decoding time to be $2^{(2mL+B(B-4m))/(B-2m)}$. So the following condition ensures $\kappa$-bit security.
$L>2m\mbox{ and either }B<2m\mbox{ or }\kappa<\frac{2mL+B(B-4m)}{B-2m}.$ (28)
@@
Evaluation against $\mathcal{S}(L,m)$. Again we have $\varepsilon=2^{-m-1}$. Setting $N=2^{B}$, we obtain k=L+2w(m+1)wBk=L+2w(m+1)-wB. Setting $T$ to be the time complexity of the decoding step, we obtain $\log_{2}T=L+2w(2m+3)-wB$. Since $L\geq 2\kappa$, ensuring $2(2m+3)\geq B$ is (more than) sufficient to ensure $\kappa$-bit security, irrespective of the values of $w$ and $k$. We record this condition as follows.
@@
Evaluation against $\mathcal{S}(L,m)$. The size of the support of the Walsh transform of ${\sf MM}_{2m+1}$ is equal to the size of the support of the Walsh transform of ${\sf MM}_{2m}$. Since ${\sf MM}_{2m}$ is bent, it follows that the size of the support of ${\sf MM}_{2m}$ is $2^{2m}$. So $F=2^{2m}$. Setting $N=2^{B}$, we obtain k=LwBk=L-wB. Substituting $k$ in the expression for $T$, we obtain $\log_{2}T=L+w(2m-B)$. Since $L\geq 2\kappa$ and $w\geq 2$, we have $\log_{2}T=L+w(2m-B)\geq 2\kappa+2(2m-B)$. So if $2\kappa+2(2m-B)>\kappa$, or equivalently, $\kappa+4m>2B$, then $\kappa$-bit security is achieved against the attack. We record this condition as follows.
@@
For a filtering function $f$ having ${\sf AI}(f)=a$, an algebraic attack requires about $N=\sum_{i=0}^{a}{L\choose i}$ keystream bits and has time complexity $T$ to be about $N^{\omega}$, where $\omega$ is the exponent of matrix multiplication (see Section 3.1.5 of [14]). Following [23] we take the value of $\omega$ to be given by Strassen’s algorithm, i.e. $\omega=\log_{2}7\approx 2.8$. For our proposal, the filtering function is $f_{2m+1}$ whose algebraic resistance is the same as that of $({\sf Maj},{\sf rev})\mbox{-}{\sf MM}_{2m+1}$. Let $a={\sf AI}(f_{2m+1})={\sf AI}(({\sf Maj},{\sf rev})\mbox{-}{\sf MM}_{2m+1})$. From Proposition 4, we assume that $a=\lceil m/2\rceil$, i.e. the actual algebraic immunity of $f_{2m+1}$ is equal to the lower bound on the algebraic immunity. Let
@@
For any $(e,d)$ for which there are non-zero functions $g$ and $h$ of degrees $e$ and $d$ respectively such that $gf_{2m+1}=h$, the Berlekamp-Massey step in a fast algebraic attack on the filter generator model has time complexity $\mathcal{O}(ED\log D)$, where $E=\sum_{i=0}^{e}{L\choose i}$ and $D=\sum_{i=0}^{d}{L\choose i}$ (see [31] and Section 3.1.5 of [14]). This complexity dominates the overall time complexity of a fast algebraic attack. Ignoring the logarithm term, we take $T$ to be equal to $ED$. The number $N$ of keystream bits required is about $2E$. Since $h$ is non-zero, the maximum value of $e$ is one less than the algebraic immunity $a$. From Proposition 4, we again assume as above that $a=\lceil m/2\rceil$. Recall that for any Boolean function, its fast algebraic immunity is at least one more than its algebraic immunity, i.e. ${\sf FAI}(f_{2m+1})\geq a+1$. For any functions $g$ and $h$ of degrees $e$ and $d$ respectively such that $gf_{2m+1}=h$, we have $e+d\geq{\sf FAI}(f_{2m+1})\geq a+1$, i.e. $d\geq a+1-e$. Define